OpenStudy (anonymous):

what is the center and radius of the circle with equation (x+9)^2+(y+5)^2=64?

OpenStudy (mathmale):

Please compare your equation (x+9)^2+(y+5)^2=64 to the general formula of a circle with center at (h,k) and radius r: (x-h)^2 + (y-k)^2 = r^2 Think carefully as you make these comparisons. Then type in what you think are the correct values for the parameters h, k and r.